OpenStudy (anonymous):

What is the value of the normal force if the coefficient of kinetic friction is 0.22 and the kinetic frictional force is 40 newtons?

OpenStudy (anonymous):

$$F_{f} = \mu \cdot F_{n}$$ $$40N = 0.22 \cdot F_{n}$$ $$\frac{40N}{0.22} = F_{n}$$ $$181.81 = F_{n}$$

OpenStudy (anonymous):

181.81N